Find the exact value of each expression

sec 30◦ cos 30◦ − tan 60◦ cot 60◦


1
Expert's answer
2021-11-29T17:43:55-0500

Let us find the exact value of the expression

"\sec 30^{\circ}\cdot \cos 30^{\circ} − \tan 60^{\circ} \cdot\cot 60^{\circ}
=\frac{1}{\cos 30^{\circ}}\cos 30^{\circ}-\frac{\sin 60^{\circ} }{\cos 60^{\circ}}\frac{\cos 60^{\circ}}{\sin 60^{\circ}}=1-1=0."